OpenStudy (anonymous):

Halving the Kelvin temperature of a gas in a rigid container decreases the gas pressure by one half. A. True B. False @Somy

OpenStudy (anonymous):

i think its true

OpenStudy (somy):

well prove it i'll give you the initial formula and u play with it and see what u get k? so PV=nRT here u go make pressure as a subject and see what you get

OpenStudy (anonymous):

when you use pressure in a sequence like this, with no numbers i get kinda confuzzled

OpenStudy (somy):

hmm so u can't play with a formula when it has no numbers?

OpenStudy (anonymous):

it messes me up a tad

OpenStudy (somy):

okaaay actually u should learn to do it this way as its actually muuuuuch easier look \(\huge P\times V = n \times R \times T\) this is how you initially see it so imagine they are numbers what would you do if you want ONLY P on the left side? you would divide nRT BY V right? so \(\huge P = \frac {n \times R \times T}{V}\)

OpenStudy (somy):

so now can you try it but this time make T as subject?

OpenStudy (somy):

PV=nRT and nRT=PV is exactly same thing so do like i did

OpenStudy (anonymous):

so its false?

OpenStudy (somy):

try it you'll see

OpenStudy (somy):

look a the formula P= nRT/V if you were to halve T which means T/2 wouldn't your Pressure also decrease? just take a look at it

OpenStudy (abmon98):

as the temperature of an enclosed gas decreases, the particles move more slowly and have less kinetic energy. They strike the container walls with less force. Halving the Kelvin temperature of a gas in a rigid container decreases the gas pressure by half. P1/T1=P2/T2 P1=1 atm T1=298 Kelvin T2=149 P2=?? P2=149*1/298=0.5 so the statement is true